Growth of the Web from June, 1993 to June, 1996 by Matthew Gray

The World Wide Web has grown substantially in the the past three years. This paper quantifies this growth based on data collected by the World Wide Web Wanderer between June, 1993 and June, 1996. The Wanderer employed a comprehensive breadth first search of the web, counting sites it encountered. The number of sites discovered increased from 130 in 6/93 to over 230,000 in 6/96, and the growth for this period is conspicuously exponential.